3over square root 6. Simplify.
Found 2 solutions by stanbon, lenny460:
Answer by stanbon(75887)

3over square root 6
------
3/sqrt(6)
---
Multiply numerator and denominator by sqrt(6) to get:
[3sqrt(6)]/6
----
Reduce to get:
= (1/2)sqrt(6)
======================
Cheers,
Stan H.
==============

Answer by lenny460(1073)

3over square root 6. Simplify.

Multiply both the numerator and the
denominator by square root 6.
We therefore get:
3 square root 6 over 6
3 / 6 = 1/2
Please note: Multiplying the denominator
square root 6 by square root 6 we get 6.
Answer:
(1/2)square root 6



Lennox Obuong
Algebra Tutor